OpenStudy (17bohtal):

You roll a standard number cube. Find P(not 4) (a) 1/6 (b) 2/3 (c) 5/6 (d) 1/3

OpenStudy (muzzack):

A?

OpenStudy (anonymous):

yeah, its A

OpenStudy (raden):

P(not 4) = 1 - P(4) = 1 - 1/6 = 5/6

OpenStudy (anonymous):

i'm trippin'